The pickling company wants to make 1400 liters of a 6% brine(salt) solution by mixing a 15% brine solution with water. How much of each should be used to accomplish this?
I thought the equation should look something like 15%B + W = 6% times 1400. But I can't get that to work.

Answer by lwsshak3(11628)

The pickling company wants to make 1400 liters of a 6% brine(salt) solution by mixing a 15% brine solution with water. How much of each should be used to accomplish this?
**
let x=liters of 15% solution to mix with water
1400-x=liters of water to mix with 15% solution
15%x+0%(1400-x)=6%(1400)
.15x=.06(1400)
x=(.06/.15)*1400=560
1400-x=840
ans:
liters of 15% solution to mix with water=560
liters of water to mix with 15% solution=840